OpenStudy (anonymous):

A plumber charges $47.50 per plus hour plus a $65.00 service charges. Your father's firm hires him to fix some leaky pipes. Find the total charges if it takes the plumber 8 hours to complete the tasks.

OpenStudy (chaotic_butterflies):

First you should set up an equation to solve your problem. The plumber has an hourly charge for his service, so for every hour (We'll label that h) we get charged 46.50. There is a down payment of 65.00 for service charges, that means it will only be charged once. So if you put that into an equation, it should look something like 47.50h + 65 = cost. You are given that he works for 8 hours, so you should plug that into the h value. So, work it out to be 47.05 (7) + 65 = cost 329.35 + 65 = cost 394.35 = cost